OpenStudy (anonymous):

solve : x+4.25=3.5x-1.5x-0.75

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

can u tell me how everything goes into place to get the answer.

OpenStudy (anonymous):

x + 4.25 = 3.5x - 1.5x - 0.75 first step: move all like variables to the same side: x - 3.5x + 1.5x = -4.25 - 0.75 now combine: -x = -5 divide by -1 to get positive x by itself: x = 5

OpenStudy (anonymous):

step by step.

OpenStudy (anonymous):

that was step by step............